The equilibrium constant is given for one of the reactions below. Determine the value of the missing equilibrium constant.

2HD(g) ⇌ H2(g) + D2(g) Kc = 0.28
6H2(g) + 6D2(g) ⇌ 12HD(g) Kc = ?

A) 0.00048
B) 1.2
C) 1.62
D) 2075
E) 0.81

*Please explain how to do this problem! Thank you!
Answer #1

The 1st reaction is:

2 HD(g) <—> H2(g) + D2(g)

Kc1 = [H2][D2]/[HD]^2

= 0.28

The 2nd reaction is:

6 H2(g) + 6 D2(g) <—> 12 HD(g)

Kc = [HD]^12 / [H2]^6[D2]^6

= (1/Kc1)^6

= (1/0.28)^6

= (3.57)^6

= 2075

Answer: D